OpenStudy (anonymous):

In your own words, explain how to solve a quadratic equation by completing the square. Demonstrate the process with your own example. answer with 150 to 200 words

OpenStudy (anonymous):

Quadratic form ax^2+bx+c Complete the square - f(x) = -2x^2+4x+1 - separate the variables -2x^2+4x+1=0 -2x^2+4x=-1 - ensure a = 1 -2x^2+4x=-1/-2 = x^2-2x=1/2 - find b/2 -2/2 = -1 - add (b/2)^2 to both sides (-2/2)^2 = 1 x^2-2x+1 = 1/2 +1 - variable side of the equation becomes the square of a binomial (x+b/2a)^2 x^2-2x +1 = (x-1)^2 - solve for x \[\sqrt{(x-1)^2} = \pm \sqrt{3/2}\] \[x-1 = \pm \sqrt{3/2}\] \[x = \pm \sqrt{3/2} +1\] \[x = 1 + \pm \sqrt{3/2}\] \[x = 1 \sqrt{3}/\sqrt{2} * \sqrt{2}/\sqrt{2}\] \[x = 1 \pm \sqrt{6}/2 \]